OpenStudy (anonymous):

Write the equation of the line with a slope of 3/7 passing through the point (9, -2). Calculate the equation of the line between the points (-6, -1) and (2, -1). If MN is the diameter of a circle with M(-3, -4) and N(-7, -14), find the center of the circle.

OpenStudy (anonymous):

is that a 1 question

OpenStudy (anonymous):

no sir 3 seperate ?

OpenStudy (lalaly):

remainder u wna do it?

OpenStudy (anonymous):

yes

OpenStudy (lalaly):

ok:)

OpenStudy (anonymous):

laltay can u help me

OpenStudy (lalaly):

laltay? lol ok

OpenStudy (lalaly):

Write the equation of the line with a slope of 3/7 passing through the point (9, -2). \[y-(-2)=\frac{3}{7}(x-9) \] \[y=\frac{3}{7}x -\frac{41}{7}\] Calculate the equation of the line between the points (-6, -1) and (2, -1) \[slope = \frac{-1-(-1)}{2--6} = 0\]\[y--1=0(x--6) \rightarrow y=-1\]

OpenStudy (anonymous):

is that it